a sound wave is found to have a frequency of 50Hz. Calculate the time period of the wave in seconds
1
Expert's answer
2016-12-13T11:46:12-0500
The reciprocal of the time period is the frequency of the wave.
f = 1 / T
T = 1 / f
T = 1 / 50 = 0.02 s